HUBBLE SPACE TELESCOPE DAILY REPORT #1110 PERIOD COVERED: 0700 EST Daily Status Report as of 095/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 5357 (Cepheid Masses: Cycle 4 High) The High Resolution Spectrograph (Side two) was used to make ultraviolet spectral observations of binary Cepheid star HD183344 to determine its mass. Twelve observations were received and displayed in the OSS. They were reported to be satisfactory. A wavelength calibration observation was also taken. 1.2 Completed FOS/BL 5530 (FOS Location of Spectra and Aperture) The Faint Object Spectrograph (Blue Side) was used to make Y-base maps of the FOS detectors using internal lamps. The purpose is to determine the location of the spectra on the diodes. A total of nine observations were received and displayed. They all were reported to be okay. 1.3 Completed Three Sets of WF/PC-2 5561 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats) The WF/PC-2 was used to make a internal dark and a visual flat calibration images to monitor flat field stability and the stability of the analog-to-digital converters. Both observations were monitored by the OSS, and no anomalies were noted. 1.4 Completed WF/PC-2 5476 (Continuum Snapshots fo 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C227.0 as part of a continuum survey and to study the galaxy's morphology. Eight images were received and displayed. Two images were reported to have had recenterings, but looked okay. The other six images were okay. 1.5 Completed WF/PC-2 5343 (WFC Imaging of Nearby Bright Quasars) The WF/PC-2 was used to make images of the bright quasar PG1307+085 and its host galaxy. The goal is to study the morphology of this galaxy, and to see how it differs from non-QSO galaxies. They also will be looking for the relationship between the properties of the host galaxy and the luminosity of the quasar. A real-time pointing correction was successfully uplinked to the spacecraft. Three images were received and displayed. They were reported to be okay. 1.6 Completed WF/PC-2 5560 (WF/PC-2 cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and for contamination. The data remains on the science tape recorder and will be analyzed later. No anomalies were noted. 1.7 Completed FOC/96 5140 (Observations of the Narrow Line Regions in Nearby Seyfert Galaxies) The Faint Object Camera (f/96) was used to make observations at several different wavelengths of Seyfert galaxy MRK-6 which is located near the Cygnus A source. This is to study relations between the radio and optical emissions. The OSS monitored the observations, and assessed the data quality. The science data still remains on the science tape recorder. No anomalies were noted. 1.8 Completed WF/PC-2 5092 (GTO Low Latitude Parallel Program: Cycle 4) The WF/PC-2 was used to make observations taken in parallel with the prime science observations. This is part of a survey to take pictures of a large number of fields within 30 degrees of the galactic plane. The intent is to determine the luminosity function of low mass stars and the mass density in the solar neighborhood. The observations were made during LOS, and remain on the science tape recorder. They will be analyzed later. No anomalies were noted. 1.9 Completed WF/PC-2 5612 (SMOV WF/PC-2 Decontamination) The WF/PC-2 was used to perform decontamination operations. Observations of standard A+81D266, and K-Spot observations were also taken. The OSS monitored the observations, and assessed the data quality. No anomalies were noted. 1.10 Completed WF/PC-2 5415 (The Distance to the M-96 Galaxy Group: Cycle 4 Medium) The WF/PC-2 was used to make observations of Cepheid variable stars within the galaxy UGC-5889 which itself is located in the M-96 galaxy cluster. This is to help determine the distance to this galactic group, and to improve the determination of the Hubble constant. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed GHRS/2 5506 (The Deuterium Abundance in Moderate Redshift QSO Systems) The High Resolution Spectrograph (Side Two) was used to make observations of QSO 0454-220. A wavelength calibration image was also made. The observations were executed as scheduled, and monitored in the OSS. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled :18 Successful :18 2.3 Significant HSTARs: None 2.4 Operations Notes: Yesterday's reports on the solar array anomaly were partially in error and require some clarification as to the time of occurrence. The event actually happened on day 88 (Tuesday) at approximately 1738Z, but due to the benign nature of the problem it was not detected during normal operations. It was observed on day 91 (Friday) when data became available for routine detailed data analysis. The tiger team has determined that a short circuit occurred from the +E SPA hot wire to the +B SPA return wire. The +E SPA power output (about 5% of the total SA output) has been lost. The failure path within the SA is causing a misleadingly low reading of the SA section 4 current telemetry, and is causing battery 5 to receive less charge than the others. Normal science operations have been ongoing since the failure. Reconfiguration of the EPS subsystem is being contemplated for the near future to cause battery 5 to receive charge commensurate with the others. For the past 24 hours, there have been no additional problem reports from any HST subsystem. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 5299 (Highly Ionized Gas in the Warped Outer Galaxies: Cycle 4 Medium) 3.2 Complete WF/PC-2 5561 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats) 3.3 Complete WF/PC-2 5415 (The Distance to the M-96 Galaxy Group: Cycle 4 Medium) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-